OpenStudy (anonymous):

f(x)=(x^4)/4 - 2x+2 and g(x)= 2x+2. find intersection points.

OpenStudy (anonymous):

i have no idea how to isolate the x. tried moving everything to one side..

OpenStudy (dumbcow):

ok what did you get when you moved everything to 1 side

OpenStudy (anonymous):

(x^4)/4 -4x

OpenStudy (dumbcow):

good now multiply equation by 4 to get rid of fraction then factor an x out

OpenStudy (anonymous):

4((x^4)-x)=0

OpenStudy (anonymous):

(x^4)-x=0

OpenStudy (dumbcow):

no should be x^4 - 16x = 0

OpenStudy (dumbcow):

x(x^3 -16) = 0

OpenStudy (anonymous):

wait, why 16x?

OpenStudy (anonymous):

oh. ok. i see.

OpenStudy (dumbcow):

\[(\frac{x^{4}}{4} -4x) *4 = x^{4} -16x\]

OpenStudy (anonymous):

so x=0 and x=\[\sqrt[3]{16}\]

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

thank you, cow! you're not dumb.

OpenStudy (anonymous):

have a great day/night.